# CT-Mamba: A Hybrid Convolutional State Space Model for Low-Dose CT Denoising

### v2

@@ Abstract
Low-dose CT (LDCT) significantly reduces the radiation dose received by patients, however, dose reduction introduces additional noise and artifacts. Currently, denoising methods based on convolutional neural networks (CNNs) face limitations in long-range modeling capabilities, while Transformer-based denoising methods, although capable of powerful long-range modeling, suffer from high computational complexity. Furthermore, the denoised images predicted by deep learning-based techniques inevitably exhibit differences in noise distribution compared to normal-dose CT (NDCT) images, which can also impact the final image quality and diagnostic outcomes. This paper proposes CT-Mamba, a hybrid convolutional State Space Model for LDCT image denoising. The model combines the local feature extraction advantages of CNNs with Mamba’s strength in capturing long-range dependencies, enabling it to capture both local details and global context. Additionally, we introduce an innovative spatially coherent ‘Z’-shaped scanning scheme to ensure spatial continuity between adjacent pixels in the image. We design a Mamba-driven deep noise power spectrum (NPS) loss function to guide model training, ensuring that the noise texture of the denoised LDCT images closely resembles that of NDCT images, thereby enhancing overall image quality and diagnostic value. Experimental results have demonstrated that CT-Mamba performs excellently in reducing noise in LDCT images, enhancing detail preservation, and optimizing noise texture distribution, and exhibits higher statistical similarity with the radiomics features of NDCT images. The proposed CT-Mamba demonstrates outstanding performance in LDCT denoising and holds promise as a representative approach for applying the Mamba framework to LDCT denoising tasks. Our code will be made available after the paper is officially published: https://github.com/zy2219105/CT-Mamba/.
Index Terms:
@@
I Introduction
@@
Computed tomography (CT) is an essential imaging technique in clinical practice, providing crucial anatomical information that aids physicians in making appropriate medical decisions. However, frequent CT scans can significantly increase the radiation dose received by patients, particularly in radiation therapy, where multiple scans are often required to achieve precise tumor localization and minimize damage to organs at risk. Excessive radiation may harm the patient, impact organ function, increase the incidence of radiation-related diseases linked to genetic damage, and ultimately reduce the patient’s quality of life. Therefore, reducing the radiation dose of CT imaging has become a focus of attention. Low-dose CT (LDCT) is an effective method for reducing patient radiation exposure [1]. In clinical practice, to acquire LDCT images, it is common to reduce the tube current to decrease the flux of the electron beam that generates X-ray photons, and/or reduce the number of projections during CT procedures. While these methods reduce radiation dose, they also introduce additional noise and artifacts, which decrease the overall quality of the image. If these issues are not effectively addressed, they will significantly impact the application of LDCT in various clinical scenarios. How to obtain high-quality CT images that are comparable to normal-dose CT (NDCT) images and meet the clinical utilities from low-dose scanning protocols has become a long-standing and practically significant problem in the field of CT imaging.
@@
To address this issue, several LDCT imaging algorithms have been proposed, which can be roughly divided into the following three categories [2, 3]: sinogram-based preprocessing, iterative reconstruction, and image post-processing. Sinogram-based preprocessing methods directly target the raw data acquisition stage and designs specific filters for processing CT projection data obtained through low-dose X-rays. Typical methods include bilateral filtering [4], structure-adaptive filtering [5], and penalized weighted least squares [6]. These denoising methods combine physical characteristics with photon statistical characteristics, making them relatively simple and easy to implement. However, these methods rely heavily on high-quality original projection data, limiting their ability to effectively restore undersampled or missing signals, and their practical application is limited by the difficulty in obtaining complete sinogram data. Iterative reconstruction methods [7] typically reconstruct images by iteratively optimizing an objective function. It usually alternates between forward and backward projections in the projection domain and the image domain until the objective function is minimized based on the convergence criterion. However, this method typically requires access to the raw data and bears a high computational cost. In addition, the quality of image reconstruction heavily relies on the precise setting of the objective function and hyperparameters. These limitations hinder the application of iterative reconstruction methods in clinical practice.
@@
CNN-based methods effectively extract image features through convolutional layers, and have achieved remarkable success in tasks such as image restoration [12], object detection [13], and image segmentation [14]. In the field of LDCT denoising, several representative CNN-based models have been proposed [15, 16, 17, 18, 19, 20, 21]. However, CNN frameworks are limited by their lack of long-range modeling capabilities, making it difficult to capture the global contextual information in images [22]. Transformers were originally designed for natural language processing, but have now been successfully adapted for image processing, such as Vision Transformer (ViT) [23] for image recognition and SwinTransformer [24] for a general backbone in various vision tasks. By treating images as sequences of patches, Transformers has better capabilities to capture global information. However, they are computationally expensive, and this is because the self-attention mechanism scales quadratically with the input size [25]. This characteristic creates a serious bottleneck, particularly in high-resolution applications and edge computing scenarios. To overcome these limitations, researchers are working to improve these models.
@@
II Medthod
II-A Preliminary: State Space Model
@@ II-B Framework Overview
We designed a Multi-Scale Coherence Mamba architecture (MSC-Mamba) to learn the features of each frequency branch obtained by wavelet transform decomposition, as shown in Fig. 1(b). MSC-Mamba captures and fuses frequency features learned from three different scales (32×32, 16×16, and 8×8) by integrating Coherence Z-Scan State Space Block (CZSS). At lower scales, MSC-Mamba effectively captures global feature information, while at higher scales, it focuses more on local details. By combining feature learning across different scales, MSC-Mamba effectively ensures spatial coherence, providing superior denoising performance. Additionally, inspired by [35], we designed a Frequency Domain Attention Module (FDAM) based on Fourier transform to enhance the low-frequency features in the wavelet domain, as shown in Fig. 3. Each grid in the extracted Fourier spectrum contains global information on the low-frequency features, enabling efficient modeling of long-range spatial dependencies in the frequency domain. Enhanced by FDAM, richer low-frequency characteristics are provided for MSC-Mamba located in the low-frequency branch, thereby optimizing the modeling of overall structural information and improving CT-Mamba’s performance.
@@
II-C Coherence Z-Scan State Space Block
@@
II-D Deep Noise Power Spectrum Loss
@@ II-E Overall Objective
3) Feature-Level Perceptual Loss: Perceptual loss captures information at the feature level using a pre-trained deep network. This loss is based on ResNet50 [36], pre-trained on the ImageNet dataset, with its weights frozen during training. Since the pre-trained ResNet50 accepts color images as input, we replicated the grayscale CT images across three channels to fit this input requirement. we extracted feature maps after its four main stages and used L2 loss to measure the similarity between the CT-Mamba output and the ground truth image. The perceptual loss is defined as (8):
@@
III EXPERIMENT DESIGNS AND RESULTS
III-A Datasets
In this work, we used a publicly released patient dataset for the 2016 NIH-AAPM-Mayo Clinic Low-Dose CT Grand Challenge. This dataset contains paired low-dose and normal-dose abdominal images with a slice thickness of 1 mm for 10 patients. The quarter-dose LDCT images were generated by adding Poisson noise to the projection data of the NDCT images to mimic a noise level that corresponded to 25% of the full dose. We used data from 9 patients for model training, totaling 5,410 image pairs, while data from the remaining patient (L506) was used for testing model performance, totaling 526 pairs. Additionally, using the method in [37], we simulated paired low-dose and normal-dose images for 20 patients from the AMOS dataset (referred to as the Simulator dataset) for model training and validation. The low-dose images also correspond to a quarter of the normal dose. We used data from 19 patients for training, totaling 1,872 image pairs, and data from 1 patient (AMOS-0033) for testing model performance, totaling 106 pairs.
III-B Implementation details
@@ III-C1 Visual evaluation
III-C2 Quantitative analysis
In Table. I and Table. II, we present the average quantitative results for PSNR, SSIM, RMSE, and VIF for patient L506 from the Mayo dataset and patient AMOS-0033 from the Simulator dataset, respectively. A comprehensive evaluation of the PSNR, SSIM, RMSE, and VIF metrics shows that CT-Mamba model outperforms the comparison methods across all quantitative indicators.
TABLE I: The average quantitative results of different methods for patient(L506) in the Mayo test dataset.
@@ III-C2 Quantitative analysis
TABLE II: The average quantitative results of different methods for patient(AMOS-0033) in the Simulator test dataset.
@@
TABLE III: The p-values calculated using the Wilcoxon rank-sum test for each model. AP-value of under 0.05 indicates statistical difference. Pairs with statistical differences are marked by *.
@@
TABLE IV: Number of similarity ratio which are higher for each model compared to the LDCT volumes out of the 93 features.
@@
III-C3 Radiomics reserach
(1) Statistical Distribution of First-Order Features: In statistical analysis, the Wilcoxon rank-sum test is commonly used to assess whether two datasets share the same distribution. For each volume, the Wilcoxon rank-sum test was employed to evaluate the consistency of radiomics features distribution between the denoised images and the reference NDCT (p $>$ 0.05). In Table III, we show the p-values for each model across the target organs (aorta, right kidney, liver, stomach, small bowel, and urinary bladder). A higher p-value indicates greater similarity to the radiomics features distribution of the NDCT. As shown in Table III, while VM-UNet performed well for the liver and stomach, its performance for the aorta was poor, exhibiting significant differences in radiomics features distributios compared to the NDCT. In contrast, the proposed CT-Mamba model demonstrated superior performance across multiple organs, including the aorta, right kidney, and urinary bladder. Furthermore, its radiomics features distribution closely aligned with those of the NDCT for all other target organs. Notably, for the aorta, the CT-Mamba outperformed all other models.
$R_{i}=\left|\frac{V_{\text{NDCT}}-V_{i}}{V_{\text{NDCT}}}\right|,$ (10)
As shown in Table IV, the proposed CT-Mamba demonstrated superior performance across all target organs. Notably, for the small bowel, where other models performed poorly, CT-Mamba achieved the best performance, further highlighting its robustness and effectiveness. The radiomics analysis indicate that the proposed CT-Mamba effectively preserves the shape and texture features of multiple organs within the LDCT images, demonstrating its potential value for clinical applications.
III-C4 Ablation study
@@
(A2) Effectiveness of the Frequency Domain Attention Module: Excluding the FDAM module.
@@
The quantitative results of each ablation experiment are summarized in Table. V. The results indicate that the complete CT-Mamba model achieves the best performance across all metrics, further validating the significant contribution of each component to the overall performance of model.
TABLE V: Quantitative comparison of different ablation experiments for patient(L506) in the Mayo test dataset.
@@
Figure 10: (a) Slice of patient L506 from the Mayo dataset shown in Figure 9, with the red dashed box indicating the uniform ROI used for NPS analysis. (b) Radial NPS comparison of this ROI, showing the difference in noise distribution between the model without deep NPS loss (A1) and the complete model with deep NPS loss (M).
IV CONCLUSIONS

### v3

@@ Abstract
Low-dose CT (LDCT) significantly reduces the radiation dose received by patients, however, dose reduction introduces additional noise and artifacts. Currently, denoising methods based on convolutional neural networks (CNNs) face limitations in long-range modeling capabilities, while Transformer-based denoising methods, although capable of powerful long-range modeling, suffer from high computational complexity. Furthermore, the denoised images predicted by deep learning-based techniques inevitably exhibit differences in noise distribution compared to normal-dose CT (NDCT) images, which can also impact the final image quality and diagnostic outcomes. This paper proposes CT-Mamba, a hybrid convolutional State Space Model for LDCT image denoising. The model combines the local feature extraction advantages of CNNs with Mamba’s strength in capturing long-range dependencies, enabling it to capture both local details and global context. Additionally, we introduce an innovative spatially coherent ‘Z’-shaped scanning scheme to ensure spatial continuity between adjacent pixels in the image. We design a Mamba-driven deep noise power spectrum (NPS) loss function to guide model training, ensuring that the noise texture of the denoised LDCT images closely resembles that of NDCT images, thereby enhancing overall image quality and diagnostic value. Experimental results have demonstrated that CT-Mamba performs excellently in reducing noise in LDCT images, enhancing detail preservation, and optimizing noise texture distribution, and exhibits higher statistical similarity with the radiomics features of NDCT images. The proposed CT-Mamba demonstrates outstanding performance in LDCT denoising and holds promise as a representative approach for applying the Mamba framework to LDCT denoising tasks.
keywords:
@@
[…]
[label1]organization=School of Physics, Beihang University, city=Beijing, country=China
[…]
[label2]organization=Emergency Department of the Third Medical Center, Chinese PLA General Hospital,city=Beijing, country=China
[…]
[label3]organization=Hangzhou International Innovation Institute, Beihang University,city=Hangzhou, country=China
[…]
[label4]organization=Tianmushan Laboratory,city=Hangzhou, country=China
1 Introduction
11footnotetext: Linxuan Li and Wenjia Wei are co-first authors.22footnotetext: Github link: https://github.com/zy2219105/CT-Mamba33footnotetext: Contact Email: zy2219105@buaa.edu.cn, weiwenjia@buaa.edu.cn
@@ 1 Introduction
Computed tomography (CT) is an essential imaging technique in clinical practice, providing crucial anatomical information that aids physicians in making appropriate medical decisions [49]. However, frequent CT scans can significantly increase the radiation dose received by patients, particularly in radiation therapy, where multiple scans are often required to achieve precise tumor localization and minimize damage to organs at risk. Excessive radiation may harm the patient, impact organ function, increase the incidence of radiation-related diseases linked to genetic damage, and ultimately reduce the patient’s quality of life. Therefore, reducing the radiation dose of CT imaging has become a focus of attention. Low-dose CT (LDCT) is an effective method for reducing patient radiation exposure [2]. In clinical practice, to acquire LDCT images, it is common to reduce the tube current to decrease the flux of the electron beam that generates X-ray photons, and/or reduce the number of projections during CT procedures. While these methods reduce radiation dose, they also introduce additional noise and artifacts, which decrease the overall quality of the image. If these issues are not effectively addressed, they will significantly impact the application of LDCT in various clinical scenarios. How to obtain high-quality CT image‘s that are comparable to normal-dose CT (NDCT) images and meet the clinical utilities from low-dose scanning protocols has become a long-standing and practically significant problem in the field of CT imaging.
@@
To address this issue, several LDCT imaging algorithms have been proposed, which can be roughly divided into the following three categories [48, 23, 51]: sinogram-based preprocessing, iterative reconstruction, and image post-processing. Sinogram-based preprocessing methods directly target the raw data acquisition stage and designs specific filters for processing CT projection data obtained through low-dose X-rays. Typical methods include bilateral filtering [29], structure-adaptive filtering [1], and penalized weighted least squares [39]. These denoising methods combine physical characteristics with photon statistical characteristics, making them relatively simple and easy to implement. However, these methods rely heavily on high-quality original projection data, limiting their ability to effectively restore undersampled or missing signals, and their practical application is limited by the difficulty in obtaining complete sinogram data. Iterative reconstruction methods [4] typically reconstruct images by iteratively optimizing an objective function. It usually alternates between forward and backward projections in the projection domain and the image domain until the objective function is minimized based on the convergence criterion. However, this method typically requires access to the raw data and bears a high computational cost. In addition, the quality of image reconstruction heavily relies on the precise setting of the objective function and hyperparameters. These limitations hinder the application of iterative reconstruction methods in clinical practice.
@@
CNN-based methods effectively extract image features through convolutional layers, and have achieved remarkable success in tasks such as image restoration [7], object detection [34], and image segmentation [35]. In the field of LDCT denoising, several representative CNN-based models have been proposed [32, 24, 11, 21, 31, 50, 20]. However, CNN frameworks are limited by their lack of long-range modeling capabilities, making it difficult to capture the global contextual information in images [47, 52]. Transformers were originally designed for natural language processing, but have now been successfully adapted for image processing, such as Vision Transformer (ViT) [6] for image recognition and SwinTransformer [27] for a general backbone in various vision tasks. By treating images as sequences of patches, Transformers has better capabilities to capture global information. However, they are computationally expensive, and this is because the self-attention mechanism scales quadratically with the input size [28]. This characteristic creates a serious bottleneck, particularly in high-resolution applications and edge computing scenarios. To overcome these limitations, researchers are working to improve these models.
@@
2 Methodology
2.1 Preliminary: State Space Model
@@ 2.2 Framework Overview
We designed a Multi-Scale Coherence Mamba architecture (MSC-Mamba) to learn the features of each frequency branch obtained by wavelet transform decomposition, as shown in Fig. 1(b). MSC-Mamba captures and fuses frequency features learned from three different scales ($H\times W\times C$, $\frac{H}{2}\times\frac{W}{2}\times 4C$, and $\frac{H}{4}\times\frac{W}{4}\times 16C$) by integrating Coherence Z-Scan State Space Block (CZSS). At lower scales, MSC-Mamba effectively captures global feature information, while at higher scales, it focuses more on local details. By combining feature learning across different scales, MSC-Mamba effectively ensures spatial coherence, providing superior denoising performance. Additionally, inspired by [14], we designed a Frequency Domain Attention Module (FDAM) based on Fourier transform to enhance the low-frequency features in the wavelet domain, as shown in Fig. 3. Each grid in the extracted Fourier spectrum contains global information on the low-frequency features, enabling efficient modeling of long-range spatial dependencies in the frequency domain. Enhanced by FDAM, richer low-frequency characteristics are provided for MSC-Mamba located in the low-frequency branch, thereby optimizing the modeling of overall structural information and improving CT-Mamba’s performance.
@@
2.3 Coherence Z-Scan State Space Block
@@
2.4 Deep Noise Power Spectrum Loss
@@ 2.5 Overall Objective
3) Feature-Level Perceptual Loss: Perceptual loss captures information at the feature level using a pre-trained deep network. This loss is based on ResNet50 [15], pre-trained on the ImageNet dataset, with its weights frozen during training. Since the pre-trained ResNet50 accepts color images as input, we replicated the grayscale CT images across three channels to fit this input requirement. This approach has been validated in LDCT tasks [46]. we extracted feature maps after its four main stages and used L2 loss to measure the similarity between the CT-Mamba output and the ground truth image. The perceptual loss is defined as (8):
@@
3 EXPERIMENT DESIGNS AND RESULTS
3.1 Datasets
In this work, we used a publicly released patient dataset for the 2016 NIH-AAPM-Mayo Clinic Low-Dose CT Grand Challenge [30]. This dataset contains paired low-dose and normal-dose abdominal images with a slice thickness of 1 mm for 10 patients. The quarter-dose LDCT images were generated by adding Poisson noise to the projection data of the NDCT images to mimic a noise level that corresponded to 25% of the full dose. We used data from 9 patients for model training, totaling 5,410 image pairs, while data from the remaining patient (L506) was used for testing model performance, totaling 526 pairs. Additionally, using the method in [40], we simulated paired low-dose and normal-dose images for 20 patients from the AMOS dataset (referred to as the Simulator dataset) for model training and validation [18]. The low-dose images also correspond to a quarter of the normal dose. We used data from 19 patients for training, totaling 1,872 image pairs, and data from 1 patient (AMOS-0033) for testing model performance, totaling 106 pairs.
3.2 Implementation details
@@ 3.3.1 Visual evaluation
Table 1: The average quantitative results of different methods for patient(L506) in the Mayo test dataset.
@@
Table 2: The average quantitative results of different methods for patient(AMOS-0033) in the Simulator test dataset.
@@
3.3.2 Quantitative analysis
In Table. 1 and Table. 2, we present the average quantitative results for PSNR, SSIM, RMSE, and VIF for patient L506 from the Mayo dataset and patient AMOS-0033 from the Simulator dataset, respectively. A comprehensive evaluation of the PSNR, SSIM, RMSE, and VIF metrics shows that CT-Mamba model outperforms the comparison methods across all quantitative indicators.
3.3.3 Radiomics reserach
(1) Statistical Distribution of First-Order Features: In statistical analysis, the Wilcoxon rank-sum test is commonly used to assess whether two datasets share the same distribution. For each volume, the Wilcoxon rank-sum test was employed to evaluate the consistency of radiomics features distribution between the denoised images and the reference NDCT (p $>$ 0.05). In Table 3, we show the p-values for each model across the target organs (aorta, right kidney, liver, stomach, small bowel, and urinary bladder). A higher p-value indicates greater similarity to the radiomics features distribution of the NDCT. As shown in Table 3, while VM-UNet performed well for the liver and stomach, its performance for the aorta was poor, exhibiting significant differences in radiomics features distributios compared to the NDCT. In contrast, the proposed CT-Mamba model demonstrated superior performance across multiple organs, including the aorta, right kidney, and urinary bladder. Furthermore, its radiomics features distribution closely aligned with those of the NDCT for all other target organs. Notably, for the aorta, the CT-Mamba outperformed all other models.
$R_{i}=\left|\frac{V_{\text{NDCT}}-V_{i}}{V_{\text{NDCT}}}\right|,$ (10)
As shown in Table 4, the proposed CT-Mamba demonstrated superior performance across all target organs. Notably, for the small bowel, where other models performed poorly, CT-Mamba achieved the best performance, further highlighting its robustness and effectiveness. The radiomics analysis indicate that the proposed CT-Mamba effectively preserves the shape and texture features of multiple organs within the LDCT images, demonstrating its potential value for clinical applications.
Table 3: The p-values calculated using the Wilcoxon rank-sum test for each model. AP-value of under 0.05 indicates statistical difference. Pairs with statistical differences are marked by *.
@@ 3.3.3 Radiomics reserach
Table 4: Number of similarity ratio which are higher for each model compared to the LDCT volumes out of the 93 features.
@@
3.3.4 Ablation study
@@
The quantitative results of each ablation experiment are summarized in Table. 5. The results indicate that the complete CT-Mamba model achieves the best performance across all metrics, further validating the significant contribution of each component to the overall performance of model.
Table 5: Quantitative comparison of different ablation experiments for patient(L506) in the Mayo test dataset.
@@
Figure 10: (a) Slice of patient L506 from the Mayo dataset shown in Fig. 8, with the red dashed box indicating the uniform ROI used for NPS analysis. (b) Radial NPS comparison of this ROI, showing the difference in noise distribution between the model without deep NPS loss (A1) and the complete model with deep NPS loss (M).
4 CONCLUSIONS
@@
CRediT authorship contribution statement
Linxuan Li: Writing - Original Draft, Methodology, Software, Conceptualization, Investigation. Wenjia Wei: Writing - Original Draft, Software, Methodology. Luyao Yang: Writing - Original Draft, Supervision, Validation. Wenwen Zhang: Visualization. Jiashu Dong: Data curation. Yahua Liu: Writing – review and editing. Wei Zhao: Writing – review and editing, Funding acquisition, Methodology, Conceptualization, Investigation.
Acknowledgments
This work was supported in part by the Natural Science Foundation of Zhejiang Province (No. LZ23A050002), Natural Science Foundation of Beijing (No. L246073), Natural Science Foundation of China (No. 12175012), the “111 center” (No. B20065), and the Fundamental Research Funds for the Central Universities, China.